# Doubly-Robust LLM-as-a-Judge: Externally Valid Estimation with Imperfect Personas

### v1

@@ 4 Experiments
  • Synthetic: All nuisance functions and target parameters are fully known (see Appendix E).
  • Semi-Synthetic: We sample 1000 real user conversations from PRISM (Kirk et al., 2024) and obtain the “ground truth” target parameter $\theta_{t}$ by treating ratings returned by an LLM-as-a-judge as human ratings ($Y$). We sample 50 such LLM ratings per item. Here, true nuisance functions are unknown. We sample persona ratings $\widehat{Y}$ by adding controlled error to the LLM-as-a-judge ratings (see $\S$ 4.1). This dataset instructs raters to assess the “helpfulness” of outputs on a 1-100 scale.
  • Real-World: We sample 300 real user conversations, rater characteristics (e.g., age, race), and human ratings ($Y$) from DICES (Aroyo et al., 2023), resulting in 300 conversations each with 25 human ratings each. We then sample persona ratings $\widehat{Y}$ by (i) adding controlled error to human ratings (see $\S$ 4.1) and (ii) via an LLM-as-a-judge with persona-based prompting. This dataset instructs raters to assess the “harmfulness” of outputs on a 1-4 scale.
In addition to providing a foundation for validating our doubly-robust estimator, the PSF offers a resource for the community to test future evaluation approaches under evaluation sampling bias.
@@ 4.1 Dataset Generation Procedure
We now describe how semi-synthetic (PRISM) and real-world (DICES) datasets are generated in the PSF. Further setup details, including prompts used for synthetic dataset generation, are reported in Appendix E.
@@
Source and Target Populations. In the semi-synthetic dataset (PRISM), the source population consists of conversations where users are prompted to engage in controversial topics, while the target population consists of conversations with no guided prompts. In the real-world dataset (DICES), the source population contains 350 single-turn conversations flagged by safety experts as containing a single harm type (e.g., misinformation, legal), while the target contains more complex conversations rated as containing multiple types of harm. In both cases, we model each sample as a single user–system exchange extracted from a multi-turn dialogue. We embed the input–output pair from each exchange into a low-dimensional space by first applying an embedding model (MiniLM-L6-v2) then projecting to 15 dimensions via UMAP (Becht et al., 2019). [^4] We selected 15 dimensions to ensure embeddings retained some predictive signal for ratings and source/target membership while keeping dimensionality low; results remained stable for $\geq 12$ dimensions. We also vary the demographic composition of raters across populations. In both PRISM and DICES, we define the source distribution $P_{s}(X)$ using marginal probabilities of rater characteristics reported in DICES, and target distribution $P_{t}(X)$ using population statistics released in the U.S. Census Bureau’s 2022 Annual Social and Economic Supplement (Guzman & Kollar, 2023).
@@
Covariate Shift. To vary the magnitude of covariate shift, we control the mixture between the source and target populations. We vary the content characteristics by controlling the proportion $\zeta\in[0,1]$ of target items contained within the source sample (sub-sampling from the full data to ensure that source and target sample sizes remain fixed). Additionally, we vary the rater distributions by taking the convex combination between all groups in each demographic stratum with normalization. The magnitude of the resulting covariate shift between samples is then given by the Sinkhorn Distance $\Delta\!\left(W^{s},W^{t}\right)$ (Feydy et al., 2019), where recall that $W=(X,V)$ and $V$ denotes the embedded content characteristics (MiniLM-L6-v2 + UMAP). We report the Sinkhorn distance normalized by subtracting the baseline case where there is no covariate shift for semi-synthetic and real-world settings, as it is inevitable that there will be variation in text embeddings despite sampling i.i.d. from pre-defined categories (e.g., harm types) within a population. This measure captures covariate shift resulting from content characteristics and demographic attributes in parallel.
@@
Selection Bias. We model rater attrition—when raters fail to provide a rating due to failed attention checks or task abandonment—by varying the probability that each item is rated. In PRISM, we prompt the LLM to output both a rating and a non-response “refusal” flag. In DICES, we use rater self-assessments of task understanding to assign attrition scores (see Appendix E). We then transform attrition scores into dropout probabilities using a Beta CDF with $\alpha=3$ (increasing $\beta$ increases selection bias). We censor ratings according to these probabilities while retaining the “true” rating. We quantify the magnitude of selection bias via the dropout rate, i.e., the probability that a rater fails to rate an item. Crucially, the dropout rates we simulate mirror those observed in practice. In DICES, 19 of 123 raters (15.4%) were excluded due to failed attention checks, while in PRISM, 104 of 1500 raters (6.9%) failed to provide ratings after completing the background survey. As we show in our results, existing methods (e.g., RePPI (Ji et al., 2025)) exhibit severe miscoverage at these empirically observed dropout rates. This underscores the importance of correcting for selection bias.
@@
Persona Quality. To systematically manipulate the quality of persona ratings, we perturb human ratings with controlled error. This error perturbation has (i) a bias parameter $\eta\in[-1,1]$, which induces a systematic shift, and (ii) a correlation parameter $\rho\in[-1,1]$, which parametrizes the Pearson correlation with human ratings. We construct the persona rating via a Cholesky-based procedure that transforms independent Gaussian noise to achieve the target correlation $\rho$:
$\hat{Y}=\text{clip}(\rho\cdot Y+\sqrt{1-\rho^{2}}\cdot Z\sigma_{Y}+\eta(y_{\max}-y_{\min}),y_{\min},y_{\max}),$
where $y_{\min},y_{\max}$ define the interval of the rating scale, and $Z$ represents independent Gaussian noise. To verify that our findings are robust to any artifacts of this perturbation process, we complement this error perturbation approach with persona ratings sampled from real LLMs.
@@ 4.3 Results
Finding 2: DR (Riesz) yields improved coverage and lower bias (MAE) than DR (Classical). Across levels of covariate shift, selection bias, and persona quality, we observe DR (Riesz) (Fig. 3; solid lines) obtains improved estimates compared to DR (Classical) (Fig. 3; dashed lines). While this behavior also appears in Synthetic (Fig. 3; left column), the gap between DR (Classical) and DR (Riesz) is especially pronounced when learning nuisance functions on embeddings of high-dimensional text (Fig. 3; PRISM, DICES). This illustrates the importance of directly estimating the re-weighting term $\alpha_{0}(W,C)$ (Equation. (4)) rather than learning $\omega_{0}(W)$ and $\pi_{0}(W)$ separately.
Finding 3: DR (Riesz) makes better use of persona ratings than baseline estimators. Several of our baselines — RePPI, PAR, and Persona-Based — use persona ratings to compute estimates. However, across levels of persona quality (Fig 3; top row), DR (Riesz) produces higher quality estimates than these baselines (with valid coverage for $\rho\geq.65$ on both PRISM and DICES). Fig. 5 extends this analysis to persona ratings obtained from real LLMs on DICES. For all LLMs, we observe that coverage of DR (Riesz) (solid) is markedly higher than that of RePPI (dashed). Further substantiating our systematic perturbation study Fig. 3 (top row), we observe that real LLMs that exhibit a higher correlation with human ratings (e.g., GPT-5; $\rho=0.43$) yield improved coverage over those with lower correlation (e.g., GPT-4o Mini; $\rho=0.23$). Furthermore, despite having lower correlation coefficients, we observe several models achieving comparable coverage to our artificially perturbed persona ratings (Fig. 3; $\rho=0.6$). Taken together, these findings illustrate that persona ratings from real LLMs-as-judges can be used to improve downstream estimates under evaluation sampling bias.
@@ 5 Related Work
Frameworks for Sample Efficient Estimation. Other works propose methods for improving statistical inference when data is scarce but ML predictions are abundant. Prediction-Powered Inference (PPI) and its computationally efficient variant PPI++ use ML predictions to tighten confidence intervals through a “rectifier term” that corrects for bias in ML predictions (Angelopoulos et al., 2023a; Chatzi et al., 2024; Fisch et al., 2024; Angelopoulos et al., 2023b). Ji et al. (2025) show PPI++ to be a special case of M-estimation with surrogate outcomes, a classical problem in causal inference (Robins et al., 1994; Robins & Rotnitzky, 1995; Tsiatis, 2006), and in turn propose recalibrated PPI (or RePPI) to offer more efficient estimation. However, these approaches fail to give valid coverage under evaluation sampling bias. We develop a doubly-robust estimator (Bang & Robins, 2005; Chernozhukov et al., 2018; 2023) that can handle covariate shift and selection bias simultaneously while making use of surrogate predictions/persona ratings. We also use “Riesz losses” (Chernozhukov et al., 2023; 2022a; 2022b) to estimate complicated nuisance parameters using generic ML learners.
@@ 7 Reproducibility Statement
We take several steps to ensure the reproducibility of our work. First, we document all theoretical assumptions required by our framework and provide complete proofs in Appendix B. Second, we provide thorough documentation of our experiment design, hyperparameters, and datasets required to reproduce our empirical results in Appendix E. Finally, we provide code necessary to reproduce our analysis in the supplementary material. We plan to release all code and datasets in our Persona Simulation Framework along with the paper so that the broader community can build upon our framework.
@@ A.1 Threats to the External Validity of Generative AI Evaluations
In the quantitative social sciences, external validity describes the extent to which findings from a study generalize to different populations, settings, and times (Findley et al., 2021). Threats to external validity have long been studied in survey research. For example, Levay et al. (2016) found notable discrepancies between the demographic composition of convenience samples obtained via Amazon Mechanical Turk (MTurk) versus nationally representative American National Election Study (ANES) samples. While Mullinix et al. (2015) observe that study findings often remain robust to such discrepancies, (Zhou & Fishbach, 2016) demonstrate that differential non-compliance — a form of selection bias in which participants drop out from studies non-randomly — can have substantive effects on studies’ results. Myriad factors contribute to this selection bias, such as participants’ motivation and language skills (Goodman & Paolacci, 2017).
More recently, concerns have emerged surrounding the external validity of evaluations obtained from general purpose benchmarks (e.g., MMLU, BigBench) and leaderboards (e.g., Chatbot Arena) (Ibrahim et al., 2024; Ouyang et al., 2023; Liao & Xiao, 2023). As with survey research, GenAI performance measurements can be subject to covariate shift when the distribution of system outputs or human raters differs between a lab-based evaluation and target deployment context (Saad-Falcon et al., 2023; Leemann et al., 2024; Kirk et al., 2024). Likewise, differential non-compliance can occur when raters in online rating platforms drop-out due to failed quality checks (e.g., due to poor English language proficiency) (Hsueh et al., 2009). Such selection bias can confound results if common factors (e.g., English language proficiency) affect rater drop-out and their ratings. Selection bias can also arise if some raters are more likely to voluntarily assign ratings than others — e.g., when busy physicians rate complex system outputs less frequently than more available medical students. While growing work has highlighted external validity as an important desideratum for evaluations (Ibrahim et al., 2024; Ouyang et al., 2023; Liao & Xiao, 2023), to our knowledge, no existing statistical frameworks simultaneously address covariate shift, selection bias, and high-dimensional model outputs while leveraging imperfect automated ratings.
@@ A.3 GenAI Systems as Human Surrogates in Social Science Studies
While our work foregrounds GenAI evaluation challenges, it bears conceptual and methodological similarity to work investigating GenAI systems as surrogates for human subjects in social science studies. Inline with the turn towards crowdworkers as a low-cost surrogate for target study populations ($\S$ A.1), this growing line of work introduces GenAI systems as a surrogate for more costly human subjects (Argyle et al., 2023). Notably, such work often targets the very same statistical parameters recovered by our general M-estimation framework (Table LABEL:tab:parameters). For example, let $v$ denote an item in an opinion poll (e.g., “do you believe in the right to bear arms?”), let $X$ denote rater characteristics (e.g., locale and demographics, per (Santurkar et al., 2023)), and let $Y$ represent a binary response (endorse/not endorse) to the survey item. The parameter
@@
Given the overlap between our motivating application and social science studies, we also discuss methods advancing the principled adoption of GenAI systems as surrogate data in social science research (Broska et al., 2024; Egami et al., 2024; 2023). These works view surrogate data as a flawed (Bisbee et al., 2024; Park et al., 2024; Takemoto, 2024; Abdurahman et al., 2024) but potentially informative source of information for statistical inference in social science studies. For instance, Broska et al. (2024) leverage prediction powered inference (Angelopoulos et al., 2023a) to correct for bias in surrogate data. However, as discussed in $\S$ A.4, PPI is vulnerable to covariate shift and selection bias between source and target study populations. Most related to our work, (Egami et al., 2024; 2023) introduce a doubly robust estimation approach that generalizes PPI by applying a bias-correction to the underlying moment function (as opposed to the outcome variable). Critically, however, this approach takes a design-based sampling procedure, which assumes that the probability of labeling a sample is known by the researcher in advance. This precludes the more general setting we study in our work, in which the reweighting function is unknown in advance. [^5] As noted in (Egami et al., 2024; 2023), this design-based sampling approach is well-motivated when the full corpus of documents to be annotated and corresponding sampling probabilities are known in advance. However, this assumption is violated in our setting, in which the true source/target distribution weights are unknown. As a result, the framework proposed by (Egami et al., 2024; 2023) is not directly applicable in our motivating setting with evaluation sampling bias. Having discussed both GenAI and social science applications of our framework, we now turn to the underlying statistical methodology we advance in this work.

### v2

@@ 4 Experiments
  • Fully Synthetic: All nuisance functions and target parameters are fully known (see Appendix E).
  • Semi-Synthetic PRISM: We sample 1000 real user conversations from PRISM (Kirk et al., 2024) and obtain the “ground truth” target parameter $\theta_{t}$ by treating ratings sampled from an LLM-as-a-judge as human ratings ($Y$). We sample 50 such LLM ratings per item. Unlike the fully synthetic setting, true nuisance functions are unknown. We sample persona ratings $\widehat{Y}$ by adding controlled error to the LLM-as-a-judge ratings (see $\S$ 4.1). This task instructs raters to rate the “helpfulness” of outputs on a 1-100 scale.
  • Semi-Synthetic DICES: We sample real user conversations, rater characteristics (e.g., age, race), and human ratings ($Y$) from DICES (Aroyo et al., 2023), resulting in 300 conversations with 25 human ratings each. We then sample persona ratings $\widehat{Y}$ by (i) adding controlled error to human ratings (see $\S$ 4.1) and (ii) via an LLM-as-a-judge with persona-based prompting. This dataset instructs raters to assess the “harmfulness” of outputs on a 1-4 scale.
Going forward, we refer to these three datasets as Synthetic, PRISM, and DICES, respectively. In addition to providing a foundation for validating our doubly-robust estimator, the PSF offers a resource for the community to test future evaluation approaches under evaluation sampling bias.
@@ 4.1 Dataset Generation Procedure
We now describe how both semi-synthetic datasets are generated in the PSF. Further setup details, including prompts used for synthetic dataset generation, are reported in Appendix E.
@@
Source and Target Populations. In the PRISM experiment, the source population consists of conversations where users are prompted to engage in controversial topics, while the target population consists of conversations with no guided prompts. In the DICES experiment, the source population contains 300 single-turn conversations flagged by safety experts as containing a single harm type (e.g., misinformation, legal), while the target contains more complex conversations rated as containing multiple types of harm. In both cases, we model each sample as a single user–system exchange extracted from a multi-turn dialogue. We embed the input–output pair from each exchange into a low-dimensional space by first applying an embedding model (MiniLM-L6-v2) then projecting to 15 dimensions via UMAP (Becht et al., 2019). [^5] We selected 15 dimensions to ensure embeddings retained predictive signal for ratings and source/target membership while keeping dimensionality low; results remained stable for $\geq 12$ dimensions (Fig. 7). We also vary the demographic composition of raters across populations. In both PRISM and DICES, we define the source distribution $P_{s}(X)$ using marginal probabilities of rater characteristics reported in DICES, and target distribution $P_{t}(X)$ using population statistics released by the U.S. Census Bureau (Guzman and Kollar, 2023).
@@
Covariate Shift. To vary the magnitude of covariate shift, we control the mixture between the source and target populations. We vary the content characteristics by controlling the proportion $\zeta\in[0,1]$ of target items contained within the source sample (sub-sampling from the full data to ensure that source and target sample sizes remain fixed). Additionally, we vary the rater distributions by taking the convex combination between all groups in each demographic stratum with normalization. The magnitude of the resulting covariate shift between samples is then given by the Sinkhorn Distance $\Delta\!\left(W^{s},W^{t}\right)$ (Feydy et al., 2019), where recall that $W=(X,V)$ and $V$ denotes the embedded content characteristics (MiniLM-L6-v2 + UMAP). We report the Sinkhorn distance normalized by subtracting the baseline case where there is no covariate shift for both semi-synthetic experiments, as it is inevitable that there will be variation in text embeddings despite sampling i.i.d. from pre-defined categories (e.g., harm types) within a population. This measure captures covariate shift resulting from content characteristics and demographic attributes in parallel.
@@
Selection Bias. We model rater attrition—when raters fail to provide a rating due to failed attention checks or task abandonment—by varying the probability that each item is rated. In PRISM, we prompt the LLM to output both a rating and a non-response “refusal” flag. In DICES, we use rater self-assessments of task understanding to assign attrition scores (see Appendix E). We then transform attrition scores into dropout probabilities using a Beta CDF with $\alpha=3$ (increasing $\beta$ increases selection bias). We censor ratings according to these probabilities while retaining the “true” rating. We quantify the magnitude of selection bias via the dropout rate, i.e., the probability that a rater fails to rate an item. The dropout rates we simulate mirror those observed in practice. In DICES, 19 of 123 raters (15.4%) were excluded due to failed attention checks, while in PRISM, 104 of 1500 raters (6.9%) failed to provide ratings after completing the background survey. As we show in our results, existing methods (e.g., RePPI (Ji et al., 2025)) exhibit severe miscoverage at these empirically observed dropout rates. This underscores the importance of correcting for selection bias.
@@
Persona Quality. To manipulate the quality of persona ratings, we perturb human ratings with controlled error. This perturbation has (i) a bias parameter $\eta\in[-1,1]$, which induces a systematic shift, and (ii) a correlation parameter $\rho\in[-1,1]$, which parametrizes the Pearson correlation between human ratings and persona ratings (see Appendix E). We verify the robustness of this perturbation approach by performing additional experiments using persona ratings sampled from real LLMs.
@@ 4.3 Results
Finding 2: DR (Riesz) yields improved coverage and lower bias (MAE) than DR (Classical). Across levels of covariate shift, selection bias, and persona quality, we observe DR (Riesz) (Fig. 3; solid lines) obtains improved estimates compared to DR (Classical) (Fig. 3; dashed lines). While this behavior also appears in Synthetic (Fig. 3; left column), the gap between DR (Classical) and DR (Riesz) is especially pronounced when learning nuisance functions on embeddings of high-dimensional text (Fig. 3; PRISM, DICES). This illustrates the importance of directly estimating the re-weighting term $\alpha_{0}(W,C)$ (Eq. 4) rather than learning $\omega_{0}(W)$ and $\pi_{0}(W)$ separately.
Finding 3: DR (Riesz) makes better use of persona ratings than baseline estimators. Several of our baselines — RePPI, PAR, and Persona-Based — use persona ratings to compute estimates. However, across levels of persona quality (Fig 3; top row), DR (Riesz) produces higher quality estimates than these baselines (with valid coverage for $\rho\geq.65$ on both PRISM and DICES). Fig. 5 extends this analysis to persona ratings obtained from varying LLMs on DICES. For all LLMs, we observe that coverage of DR (Riesz) (solid) is markedly higher than that of RePPI (dashed). Further substantiating our systematic perturbation study (Fig. 3; top row), we observe that real LLMs that exhibit a higher correlation with human ratings (e.g., GPT-5; $\rho=0.43$) yield improved coverage over those with lower correlation (e.g., GPT-4o Mini; $\rho=0.23$). Furthermore, despite having lower correlation coefficients, we observe several models achieving comparable coverage to our artificially perturbed persona ratings (Fig. 3; $\rho=0.6$). Taken together, these findings illustrate that persona ratings from real LLMs-as-judges can be used to improve downstream estimates under evaluation sampling bias.
@@ 5 Related Work
Frameworks for Sample Efficient Estimation. Other works propose methods for improving statistical inference when data is scarce but ML predictions are abundant. Prediction-Powered Inference (PPI) and its computationally efficient variant PPI++ use ML predictions to tighten confidence intervals through a “rectifier term” that corrects for bias in ML predictions (Angelopoulos et al., 2023a; Chatzi et al., 2024; Fisch et al., 2024; Angelopoulos et al., 2023b). Ji et al. (2025) show PPI++ to be a special case of M-estimation with surrogate outcomes, a classical problem in causal inference (Robins et al., 1994; Robins and Rotnitzky, 1995; Tsiatis, 2006), and in turn propose recalibrated PPI (or RePPI) to offer more efficient estimation. However, these approaches fail to give valid coverage under evaluation sampling bias. We develop a doubly-robust estimator (Bang and Robins, 2005; Chernozhukov et al., 2018; 2023) that can handle covariate shift and selection bias simultaneously while making use of surrogate predictions/persona ratings. We also use “Riesz losses” (Chernozhukov et al., 2023; 2022b; 2022a) to estimate complicated nuisance parameters using generic ML learners.
@@ 7 Reproducibility Statement
We take several steps to ensure the reproducibility of our work. First, we document all theoretical assumptions required by our framework and provide complete proofs in Appendix B. Second, we provide thorough documentation of our experiment design, hyperparameters, and datasets required to reproduce our empirical results in Appendix E. Finally, we publicly release all code and data so that the broader community can build upon our framework: https://github.com/lguerdan/doubly-robust-llm-judge.
@@ 8 Contribution Statement
9 Acknowledgments
We thank members of the Fairness, Ethics, Accountability, and Transparency (FEAT) reading group at CMU for their helpful feedback on earlier versions of this work. This research was funded by the National Institute of Standards and Technology (https://ror.org/05xpvk416) and Carnegie Mellon University (https://ror.org/05x2bcf33) AI Measurement Science and Engineering Center (AIMSEC). Luke Guerdan (ORCID: 0009-0009-3566-9429) was funded by NIST through Federal Award ID Number 60NANB24D231.
@@ A.1 Threats to the External Validity of Generative AI Evaluations
In the quantitative social sciences, external validity describes the extent to which findings from a study generalize to different populations, settings, and times (Findley et al., 2021). Threats to external validity have long been studied in survey research. For example, Levay et al. (2016) found notable discrepancies between the demographic composition of convenience samples obtained via Amazon Mechanical Turk (MTurk) versus nationally representative American National Election Study (ANES) samples. While Mullinix et al. (2015) observe that study findings often remain robust to such discrepancies, (Zhou and Fishbach, 2016) demonstrate that differential non-compliance — a form of selection bias in which participants drop out from studies non-randomly — can have substantive effects on studies’ results. Myriad factors contribute to this selection bias, such as participants’ motivation and language skills (Goodman and Paolacci, 2017).
More recently, concerns have emerged surrounding the external validity of evaluations obtained from general purpose benchmarks (e.g., MMLU, BigBench) and leaderboards (e.g., Chatbot Arena) (Ibrahim et al., 2024; Ouyang et al., 2023; Liao and Xiao, 2023). As with survey research, GenAI performance measurements can be subject to covariate shift when the distribution of system outputs or human raters differs between a lab-based evaluation and target deployment context (Saad-Falcon et al., 2023; Leemann et al., 2024; Kirk et al., 2024). Likewise, differential non-compliance can occur when raters in online rating platforms drop-out due to failed quality checks (e.g., due to poor English language proficiency) (Hsueh et al., 2009). Such selection bias can confound results if common factors (e.g., English language proficiency) affect rater drop-out and their ratings. Selection bias can also arise if some raters are more likely to voluntarily assign ratings than others — e.g., when busy physicians rate complex system outputs less frequently than more available medical students. While growing work has highlighted external validity as an important desideratum for evaluations (Ibrahim et al., 2024; Ouyang et al., 2023; Liao and Xiao, 2023), to our knowledge, no existing statistical frameworks simultaneously address covariate shift, selection bias, and high-dimensional model outputs while leveraging imperfect automated ratings.
@@ A.3 GenAI Systems as Human Surrogates in Social Science Studies
While our work foregrounds GenAI evaluation challenges, it bears conceptual and methodological similarity to work investigating GenAI systems as surrogates for human subjects in social science studies. Inline with the turn towards crowdworkers as a low-cost surrogate for target study populations ($\S$ A.1), this growing line of work introduces GenAI systems as a surrogate for more costly human subjects (Argyle et al., 2023). Notably, such work often targets the very same statistical parameters recovered by our general M-estimation framework (Table 1). For example, let $v$ denote an item in an opinion poll (e.g., “do you believe in the right to bear arms?”), let $X$ denote rater characteristics (e.g., locale and demographics, per (Santurkar et al., 2023)), and let $Y$ represent a binary response (endorse/not endorse) to the survey item. The parameter
@@
Given the overlap between our motivating application and social science studies, we also discuss methods advancing the principled adoption of GenAI systems as surrogate data in social science research (Broska et al., 2024; Egami et al., 2024; 2023). These works view surrogate data as a flawed (Bisbee et al., 2024; Park et al., 2024; Takemoto, 2024; Abdurahman et al., 2024) but potentially informative source of information for statistical inference in social science studies. For instance, Broska et al. (2024) leverage prediction powered inference (Angelopoulos et al., 2023a) to correct for bias in surrogate data. However, as discussed in $\S$ A.4, PPI is vulnerable to covariate shift and selection bias between source and target study populations. Most related to our work, (Egami et al., 2024; 2023) introduce a doubly robust estimation approach that generalizes PPI by applying a bias-correction to the underlying moment function (as opposed to the outcome variable). Critically, however, this approach takes a design-based sampling procedure, which assumes that the probability of labeling a sample is known by the researcher in advance. This precludes the more general setting we study in our work, in which the reweighting function is unknown in advance. [^6] As noted in (Egami et al., 2024; 2023), this design-based sampling approach is well-motivated when the full corpus of documents to be annotated and corresponding sampling probabilities are known in advance. However, this assumption is violated in our setting, in which the true source/target distribution weights are unknown. As a result, the framework proposed by (Egami et al., 2024; 2023) is not directly applicable in our motivating setting with evaluation sampling bias. Moreover, (Nwankwo et al., 2025) propose a doubly-robust estimation approach for batch-adaptive annotation similar to the framework put forth by (Egami et al., 2024; 2023), but focus on causal quantities such as the ATE rather than M-estimation problems.
Having discussed both GenAI and social science applications of our framework, we now turn to the underlying statistical methodology we advance in this work.